# Grounded Test-Time Adaptation for LLM Agents

### v3

@@ Abstract
Large language model (LLM)-based agents struggle to generalize to novel and complex environments, such as unseen websites or new sets of functions, due to a fundamental mismatch between their pre-training and test-time conditions. This challenge stems from two distinct failure modes: a syntactic misunderstanding of environment-specific components like observation formats, and a semantic misunderstanding of state-transition dynamics, which are only revealed at test time. To address these issues, we propose two distinct and complementary strategies for adapting LLM agents by leveraging environment-specific information available during deployment. First, an online distributional adaptation method parameterizes environmental nuances by learning a lightweight adaptation vector that biases the model’s output distribution, enabling rapid alignment with an environment response format. Second, a deployment-time dynamics grounding method employs a persona-driven exploration phase to systematically probe and learn the environment’s causal dynamics before task execution, equipping the agent with a non-parametric world model. We evaluate these strategies across diverse agentic benchmarks, including function calling and web navigation. Our empirical results show the effectiveness of both strategies across all benchmarks with minimal computational cost. We find that dynamics grounding is particularly effective in complex environments where unpredictable dynamics pose a major obstacle, demonstrating a robust path toward more generalizable and capable LLM-based agents. For example, on the WebArena multi-site split, this method increases the agent’s success rate from 2% to 23%. We release our code [^1] https://github.com/r2llab/GTTA.
@@ 1 Introduction
Large language model (LLM)-based AI agents have demonstrated remarkable capabilities across diverse tasks (li_can_2023; zhou_webarena_2024; patil_berkeley_nodate; xie_osworld_2024), yet they often fail to generalize when deployed in novel agentic environments such as web navigation (zhou_webarena_2024; he_webvoyager_2024) and function calling (patil_berkeley_nodate; yao_tau-bench_2024). These failures reflect two core challenges—the systematic mismatch between the agent’s pre-trained knowledge and the specific syntax and dynamics of the deployed environment (yang_agentoccam_2025; thil_navigating_2024). While both mismathes stem from a lack of prior knowledge, we distinguish them operationally because they require different adaptation mechanisms (parametric vs. non-parametric). To illustrate these two failure patterns, we consider an LLM agent tasked with booking a flight on a previously unseen travel website:
@@
    Syntactic Mismatch: the LLM agent’s prior knowledge does not align with environment-specific information, such as observation structure (yang_if_2024; gur_real-world_2023) and the environment’s unique syntax (lei_spider_2024; chen_agent-flan_2024). This causes parsing and context-understanding issues: a pre-trained LLM agent might try click("Search") or target destination, while the new site exposes Go and dest_field, producing invalid actions and immediate failures. (yang_if_2024; gur_real-world_2023; lei_spider_2024; chen_agent-flan_2024)
@@
    Semantic mismatch: the agent lacks an accurate, environment-specific causal model of state transitions, so it cannot predict the consequences of actions. For example, the agent may expect click("Go") to show flight results but the site instead opens a date-confirmation pop-up; without knowing this transition, the agent cannot form the correct multi-step plan and fails. (forrester_counterintuitive_1971; chae_web_2025; zhang_attacking_2024)
Current approaches to this generalization gap are ill-suited for rapid adaptation in novel environments. In addition, many methods require human-annotated or LLM-annotated demonstrations (wang_agent_2024; murty_nnetnav_2025; xu_agenttrek_2024), which can be resource-intensive or rely on LLM’s prior knowledge about the environment. On the other hand, to address semantic mismatch, explicit world modeling like (chae_web_2025) involves a heavyweight pipeline of extensive data collection and fine-tuning a separate, specialized model that is computationally expensive and struggles to generalize without being retrained. Both paradigms present significant overhead, highlighting the need for more efficient strategies that can ground an agent using only the information available at test time.
To address the the two failure modes and close the generalization gap, we propose two annotation-free strategies that operate at deployment time. (1) A parametric online adapter (parametric test-time adaptation)–a lightweight module that learns a small per-interaction bias (an adaptation vector applied to late features) to quickly align an agent’s output distribution with environment-specific syntax. Observing Go and dest_field on the page, for example, lets the adapter steer action generation toward the site’s actual element names. (2) A non-parametric dynamics grounding (non-parametric test-time adaptation)—a short, persona-guided exploration phase that iteratively collects in-context rules about state transitions (e.g., clicking Go opens a date pop-up) and supplies those rules as context for downstream planning. Both methods require only test-time observations and no trajectory annotations or expensive fine-tuning.
We evaluate our methods on function-calling and web-navigation benchmarks. Both approaches yield consistent improvements: the parametric adapter gives efficient, steady gains across sites, while dynamics grounding produces particularly large improvements on environments with unfamiliar transitions. For example, on the WebArena multi-site split (zhou_webarena_2024), dynamics grounding raises GPT4.1 success from 2% to 23%. Taken together, syntactic and semantic test-time grounding substantially narrow the deployment generalization gap and offer a practical path toward more robust LLM agents.
@@ 2 Related Work
Originating in computer vision, test-time adaptation (TTA) addresses distributional shifts between training and testing data by adapting models at inference time (wang_tent_2021; niu_efficient_2022; sun_test-time_2020). Our parametric test-time adaptation builds directly on methods that update model parameters or steering vectors at test time to minimize an unsupervised objective, such as entropy minimization or self-supervised learning (wang_tent_2021; hu_slot_2025; li_inference-time_2023). Our parametric test-time adaptation aligns with steering vector research (tan_analyzing_2025; sinii_steering_2025; panickssery_steering_2024; turner_steering_2024) in its use of latent space biases to shift output distributions. However, standard steering approaches generally apply modulation vectors to shift high-level traits  (tan_analyzing_2025), such as honesty (zou_representation_2025). We distinguish our work by tailoring this mechanism to the dynamic nature of agentic tasks: our vector is treated as a temporary parameter, updated online using the environment’s response as a self-supervisory signal, and reset per episode to ensure precise, context-dependent alignment without interference across tasks. While TTA for LLMs is emerging for tasks like math reasoning (zuo_ttrl_2025) and few-shot learning (akyurek_surprising_2025), its application to the unique, interactive, and stateful challenges of LLM agents remains underexplored. Our work is the first to systematically apply this paradigm to adapt agents to novel environment formats without supervised trajectories.
@@
Recent work has focused on incorporating environment dynamics into LLM-based planning by training a parameterized world model (fang_webevolver_2025; chae_web_2025; ding_understanding_2026; qiao_agent_2024), or using LLM to predict the next state using continuously updated environment rules (zhou_wall-e_2024). Such approaches typically require collecting a large set of observations for learning a world model, and can be computationally expensive (chae_web_2025). In contrast, our non-parametric test-time adaptation introduces a lightweight, deployment-time pipeline to generate a temporary, in-context world model from a few exploratory interactions. This aligns with a broader trend of using in-context learning for adaptation, including methods like agent workflow memory (wang_agent_2024) and retrieval-augmented generation (RAG) for planning (luo_reasoning_2023; kagaya_rap_2024; wang_learning_2024). However, our contribution lies in the systematic and automated process of generating this knowledge via persona-driven exploration, removing the need for pre-existing annotated demonstrations or a separate, trained dynamics model.
@@
Large language models (LLMs) have become the planning backbone of a wide range of agentic systems due to their strong instruction-following capabilities and generalization across diverse tasks (ouyang_training_2022; mishra_cross-task_2022; zhou_webarena_2024). Their ability to interpret complex instructions and generate coherent action plans has enabled their integration into applications such as web navigation (zhou_webarena_2024; he_webvoyager_2024; lu_weblinx_2024; deng_mind2web_2023) and function calling (patil_berkeley_nodate; yao_tau-bench_2024). Despite recent advances, LLM agents struggle to comprehend the specific syntactical input structure of different environments, such as in web navigation (yang_if_2024; gur_real-world_2023), text-to-SQL (lei_spider_2024). This is due to the fundamental mismatch between the LLM pre-training corpus and the specific agentic environment (chen_agent-flan_2024; ben-david_theory_2010).
@@ 3 Grounded Test-Time Adaptation Strategies
In this section, we detail our two complementary adaptation strategies. We first formalize the problem setup, which is designed to reflect realistic deployment scenarios and is defined by three key constraints:
@@
This setting mirrors constraints in practical benchmarks (e.g., BIRD (li_can_2023), WebArena (zhou_webarena_2024)) and necessitates methods that can leverage only unsupervised, test-time interaction. To formulate this challenge, at test-time the LLM input $\mathcal{I}$ is constructed as
@@
3.1 Parametric Test-Time Adaptation
Figure 1: Overview of parametric test-time adaptation. This figure includes an example of web navigation shopping task to illustrate how the agent adapts to new environment. (1) At the start of each episode, we initialize an adaptation vector $\delta$ as a zero vector and construct inputs to the LLM agent. (2) During task execution, the agent receives environment instructions and observations. (3) At each step, we update the adaptation vector using cross-entropy loss on the current input, and apply the adaptation vector as a bias to the LLM’s final hidden layer. This enables rapid alignment to environment-specific observation and action formats. (4) The LLM agent takes a new action with the updated vector, which shifts the model’s output distribution to better match the test-time environment.
To address syntactic mismatches that arise in novel environments, we propose parametric test-time adaptation. The objective is to parametrically adapt the model’s output distribution at test time by treating the current context (i.e. task instructions, observations, and action history) as a self-supervisory signal. This allows the model to align with local syntactic patterns, such as unique UI element labels or response formats, without requiring a deep understanding of the environment’s causal dynamics.
Our approach introduces a single, lightweight adaptation vector $\delta\in\mathbb{R}^{d}$, where $d$ is the hidden dimension of the language model. This vector acts as an additive bias to the final hidden representations $H\in\mathbb{R}^{n\times d}$ before the final projection layer. At each step of an episode, the adapted logits are computed as:
@@ 3.1 Parametric Test-Time Adaptation
where $f_{\theta,\delta}$ is the LLM parameterized by its fixed weights $\theta$ and the adaptable vector $\delta$, and $\eta$ is the learning rate. $\mathcal{I}_{1:n-1}$ denotes the input subsequence consisting of tokens 1 through n-1, and $\mathcal{I}_{2:n}$ denotes the corresponding next-token targets obtained by shifting the sequence by one position. The loss in Eq. 3 is therefore computed by predicting each token in $\mathcal{I}_{2:n}$ from its preceding context $\mathcal{I}_{1:n-1}$. This update encourages the model to internalize the syntactic features present in the immediate context by adapting its output distribution to the environment context.
@@
    Prediction: Without adaptation, the model’s logits would likely favor the action click("Search"), based on its pre-trained priors.
@@
    Correction: In the subsequent generation step, the adapted logits (Eq. 2) now assign a higher probability to the syntactically correct action, click("Go").
This process is computationally efficient, and step 2-4 is repeated at every model prediction within an episode. To prevent catastrophic forgetting and ensure that adaptation is specific to the current environment, $\delta$ is reset to zero at the beginning of each new episode. This lightweight update aligns the model’s pre-trained distribution with the specific distribution of the test-time environment.
3.2 Non-parametric Test-Time Adaptation
Figure 2: Overview of non-parametric test-time adaptation. This figure includes an example of web navigation shopping task to illustrate how the pipeline generates environment dynamics in language. (1) We synthesize diverse exploration tasks based on personas using environment descriptions. (2) An exploration agent interacts with the environment to collect interaction logs of state transitions and updates the environment rules/dynamics dynamically. (3) An LLM extracts and summarizes environment dynamics from these logs. (4) A reasoning model filters less informative rules, which are then used to augment the agent’s context during evaluation, enabling more transition-aware decision making.
LLM-based agents often fail in novel environments due to a semantic mismatch: they lack a “world model” to predict the outcomes of their actions (ha_world_2018; chae_web_2025). To address this, we propose non-parametric test-time adaptation, a strategy that proactively discovers the causal dynamics of an environment before the execution of the task. The objective is to construct a non-parametric (in language), in-context world model by performing a one-time, automated exploration. The cost of this discovery phase is amortized over all subsequent tasks performed in that environment.
@@ 3.2 Non-parametric Test-Time Adaptation
    Persona/Exploratory Goal Synthesis. An LLM is prompted with a high-level description of the environment (e.g., website purpose, API documentation) to generate a set of $N$ diverse, exploratory “personas” or ”goals”. Personas frame open-ended tasks designed to probe for non-obvious interactions. For our travel-example.com scenario, a synthesized persona might be: “As a first-time user, I want to see what happens if I try to search for a flight without selecting a date.”. We prefer personas over a less-structured exploration policy (e.g. maximizing state novelty or coverage) to guide the agent to explore complex, multi-step interactions that a naive exploration strategy might miss, thus yielding more semantically rich dynamics.
@@
    Exploration and on-the-fly Dynamics Extraction For each persona or goal, an LLM agent interacts with the environment. The agent is instructed to take novel actions to maximize its discovery of new state transitions. After each action is executed and transition into next state, we take (observation, action, new observation) to summarize into a concise human-readable rule $e$ Immediately after, the list of generated rule $\{e\}_{i=1}^{T}$ is appended to the agent to encourage taken action that’s not been taken before. In our example, the agent would execute click("Go") and log the transition from the main page to a state where a calendar pop-up is active. Then the transition on travel-example.com would be extracted as the dynamic: Executing the action click(’Go’) on the main page causes a calendar modal to appear for date selection.
@@
    Filtering and Consolidation. The set of all extracted dynamics, $E=\{e\}_{i=1}^{M}$, is passed to a reasoning model. This model filters out trivial (e.g., ”typing in a text field shows text”) or repetitive rules, producing a final, clean set of dynamics, $E_{\text{clean}}$.

### v4

# Test-Time Adaptation for LLM Agents via Environment Interaction

@@ Abstract
Large language model (LLM)-based agents struggle to generalize to novel and complex environments, such as unseen websites or new sets of functions, due to a fundamental mismatch between their pre-training and test-time conditions. This challenge stems from two distinct failure modes: a syntactic misunderstanding of environment-specific components like observation formats, and a semantic misunderstanding of state-transition dynamics, which are only revealed at test time. To address these issues, we propose two distinct strategies for adapting LLM agents by leveraging environment-specific information from interaction that is available during deployment. First, an online syntactic alignment (SA) method parameterizes environmental nuances by learning a lightweight adaptation vector that biases the model’s output distribution, enabling rapid alignment with an environment response format. Second, a deployment-time dynamics grounding (DG) method employs a persona-driven exploration phase to systematically probe and learn the environment’s causal dynamics before task execution, equipping the agent with an in-context world model. We evaluate these strategies across diverse agentic benchmarks, including function calling and web navigation. Our empirical results show the effectiveness of both strategies across all benchmarks with minimal computational cost. We find that dynamics grounding is particularly effective in complex environments where unpredictable dynamics pose a major obstacle, demonstrating a robust path toward more generalizable and capable LLM-based agents. For example, on the WebArena multi-site split, this method increases the agent’s success rate from 2% to 23%. We release our code [^1] https://github.com/r2llab/GTTA.
@@ 1 Introduction
Large language model (LLM)-based AI agents have demonstrated remarkable capabilities across diverse tasks (Li et al., 2023a; Zhou et al., 2024a; Patil et al., ; Xie et al., 2024), yet they often fail to generalize when deployed in novel agentic environments such as web navigation (Zhou et al., 2024a; He et al., 2024) and function calling (Patil et al., ; Yao et al., 2024). These failures reflect two core challenges—the systematic mismatch between the agent’s pre-trained knowledge and the specific syntax and dynamics of the deployed environment (Yang et al., 2025; Thil et al., 2024). Although both mismatches stem from a lack of prior knowledge, we distinguish them operationally because they require different adaptation mechanisms (syntactic vs. semantic). To illustrate these two failure patterns, we consider an LLM agent tasked with booking a flight on a previously unseen travel website:
@@
    Syntactic Mismatch: the LLM agent’s prior knowledge does not align with environment-specific information, such as observation structure (Yang et al., 2024; Gur et al., 2023) and the environment’s unique syntax (Lei et al., 2024; Chen et al., 2024). This causes parsing and context-understanding issues: a pre-trained LLM agent might try click(‘‘Search’’) or target destination, while the new site exposes Go and dest_field, producing invalid actions and immediate failures. (Yang et al., 2024; Gur et al., 2023; Lei et al., 2024; Chen et al., 2024)
@@
    Semantic Mismatch: the agent lacks an accurate, environment-specific causal model of state transitions, so it cannot predict the consequences of actions. For example, the agent may expect click(‘‘Go’’) to show flight results, but the site instead opens a date-confirmation pop-up; without knowing this transition, the agent cannot form the correct multi-step plan and fails. (Forrester, 1971; Chae et al., 2025; Zhang et al., 2024)
Current approaches to this generalization gap are ill-suited for rapid adaptation in novel environments. In addition, many methods require human-annotated or LLM-annotated demonstrations (Wang et al., 2024b; Murty et al., 2025; Xu et al., 2024), which can be resource-intensive or rely on LLM’s prior knowledge about the environment. On the other hand, to address semantic mismatch, explicit world modeling like (Chae et al., 2025) involves a heavyweight pipeline of extensive data collection and fine-tuning a separate, specialized model that is computationally expensive and struggles to generalize without being retrained. Both paradigms present significant overhead, highlighting the need for more efficient strategies that can ground an agent using only the information available at test time.
To address the two failure modes and close the generalization gap, we propose two annotation-free strategies that operate at deployment time. (1) Online syntactic alignment (SA)—a lightweight module that learns a small per-interaction bias (an adaptation vector applied to late features) to quickly align an agent’s output distribution with environment-specific syntax. Observing Go and dest_field on the page, for example, lets the adapter steer action generation toward the site’s actual element names. (2) Dynamics grounding (DG)—a short, persona-guided exploration phase that iteratively collects in-context rules about state transitions (e.g., clicking Go opens a date pop-up) and supplies those rules as context for downstream planning. Both methods require only test-time observations and no trajectory annotations or expensive fine-tuning.
We evaluate our methods on function-calling and web-navigation benchmarks. Both approaches yield consistent improvements: syntactic alignment gives efficient, steady gains across sites, while dynamics grounding produces particularly large improvements on environments with unfamiliar transitions. For example, on the WebArena multi-site split (Zhou et al., 2024a), dynamics grounding raises GPT-4.1 success from 2% to 23%. Taken together, syntactic and semantic test-time grounding substantially narrow the deployment generalization gap and offer a practical path toward more robust LLM agents.
@@ 2 Related Work
Originating in computer vision, test-time adaptation (TTA) addresses distributional shifts between training and testing data by adapting models at inference time (Wang et al., 2021; Niu et al., 2022; Sun et al., 2020). Our syntactic alignment builds directly on methods that update model parameters or steering vectors at test time to minimize an unsupervised objective, such as entropy minimization or self-supervised learning (Wang et al., 2021; Hu et al., 2025; Li et al., 2023b). Our syntactic alignment aligns with steering vector research (Tan et al., 2025; Sinii et al., 2025; Panickssery et al., 2024; Turner et al., 2024) in its use of latent space biases to shift output distributions. However, standard steering approaches generally apply modulation vectors to shift high-level traits  (Tan et al., 2025), such as honesty (Zou et al., 2025). We distinguish our work by tailoring this mechanism to the dynamic nature of agentic tasks: our vector is treated as a temporary parameter, updated online using the environment’s response as a self-supervisory signal, and reset per episode to ensure precise, context-dependent alignment without interference across tasks. While TTA for LLMs is emerging for tasks like math reasoning (Zuo et al., 2025) and few-shot learning (Akyürek et al., 2025), its application to the unique, interactive, and stateful challenges of LLM agents remains underexplored. Our work is the first to systematically apply this paradigm to adapt agents to novel environment formats without supervised trajectories.
@@
Recent work has focused on incorporating environment dynamics into LLM-based planning by training a parameterized world model (Fang et al., 2025; Chae et al., 2025; Ding et al., 2026; Qiao et al., 2024), or using an LLM to predict the next state using continuously updated environment rules (Zhou et al., 2024b). Such approaches typically require collecting a large set of observations for learning a world model, and can be computationally expensive (Chae et al., 2025). In contrast, our dynamics grounding introduces a lightweight, deployment-time pipeline to generate a temporary, in-context world model from a few exploratory interactions. This aligns with a broader trend of using in-context learning for adaptation, including methods like agent workflow memory (Wang et al., 2024b) and retrieval-augmented generation (RAG) for planning (Luo et al., 2023; Kagaya et al., 2024; Wang et al., 2024a). However, our contribution lies in the systematic and automated process of generating this knowledge via persona-driven exploration, removing the need for pre-existing annotated demonstrations or a separate, trained dynamics model.
@@
Large language models (LLMs) have become the planning backbone of a wide range of agentic systems due to their strong instruction-following capabilities and generalization across diverse tasks (Ouyang et al., 2022; Mishra et al., 2022; Zhou et al., 2024a). Their ability to interpret complex instructions and generate coherent action plans has enabled their integration into applications such as web navigation (Zhou et al., 2024a; He et al., 2024; Lù et al., 2024; Deng et al., 2023) and function calling (Patil et al., ; Yao et al., 2024). Despite recent advances, LLM agents struggle to comprehend the specific syntactical input structure of different environments, such as in web navigation (Yang et al., 2024; Gur et al., 2023) and text-to-SQL (Lei et al., 2024). This is due to the fundamental mismatch between the LLM pre-training corpus and the specific agentic environment (Chen et al., 2024; Ben-David et al., 2010).
@@ 3 Grounded Test-Time Adaptation Strategies
In this section, we detail our two adaptation strategies which mitigate syntactic and semantic gaps. We first formalize the problem setup, which is designed to reflect realistic deployment scenarios and is defined by three key constraints:
@@
This setting mirrors constraints in practical benchmarks (e.g., BIRD (Li et al., 2023a), WebArena (Zhou et al., 2024a)) and necessitates methods that can leverage only unsupervised, test-time interaction. To formulate this challenge, at test-time the LLM input $\mathcal{I}$ is constructed as
@@
3.1 Syntactic Alignment
Figure 1: Overview of syntactic alignment (SA). This figure includes an example of web navigation shopping task to illustrate how the agent adapts to new environment. (1) At the start of each episode, we initialize an adaptation vector $\delta$ as a zero vector and construct inputs to the LLM agent. (2) During task execution, the agent receives environment instructions and observations. (3) At each step, we update the adaptation vector using cross-entropy loss on the current input, and apply the adaptation vector as a bias to the LLM’s final hidden layer. This enables rapid alignment to environment-specific observation and action formats. (4) The LLM agent takes a new action with the updated vector, which shifts the model’s output distribution to better match the test-time environment.
To address syntactic mismatches that arise in novel environments, we propose syntactic alignment. The objective is to parametrically adapt the model’s output distribution at test time by treating the current context (i.e. task instructions, observations, and action history) as a self-supervisory signal. This allows the model to align with local syntactic patterns, such as unique UI element labels or response formats, without requiring a deep understanding of the environment’s causal dynamics.
Our approach introduces a single lightweight adaptation vector $\delta\in\mathbb{R}^{d}$, where $d$ is the hidden dimension of the language model. This vector acts as an additive bias to the final hidden representations $H\in\mathbb{R}^{n\times d}$ before the final projection layer. At each step of an episode, the adapted logits are computed as:
@@ 3.1 Syntactic Alignment
where $f_{\theta,\delta}$ is the LLM parameterized by its fixed weights $\theta$ and the adaptable vector $\delta$, and $\eta$ is the learning rate. $\mathcal{I}_{1:n-1}$ denotes the input subsequence consisting of tokens $1$ through n1n-1, and $\mathcal{I}_{2:n}$ denotes the corresponding next-token targets obtained by shifting the sequence by one position. The loss in Eq. 3 is therefore computed by predicting each token in $\mathcal{I}_{2:n}$ from its preceding context $\mathcal{I}_{1:n-1}$. This update encourages the model to internalize the syntactic features present in the immediate context by adapting its output distribution to the environment context.
@@
    Prediction: Without adaptation, the model’s logits would likely favor the action click(‘‘Search’’), based on its pre-trained priors.
@@
    Correction: In the subsequent generation step, the adapted logits (Eq. 2) now assign a higher probability to the syntactically correct action, click(‘‘Go’’).
This process is computationally efficient, and steps 2-4 are repeated at every model prediction within an episode. To prevent catastrophic forgetting and ensure that adaptation is specific to the current environment, $\delta$ is reset to zero at the beginning of each new episode. This lightweight update aligns the model’s pre-trained distribution with the specific distribution of the test-time environment.
3.2 Dynamics Grounding
Figure 2: Overview of dynamics grounding (DG). This figure includes an example of web navigation shopping task to illustrate how the pipeline generates environment dynamics in language. (1) We synthesize diverse exploration tasks based on personas using environment descriptions. (2) An exploration agent interacts with the environment to collect interaction logs of state transitions. (3) An LLM extracts and summarizes environment dynamics from these logs. (4) A reasoning model filters less informative rules, which are then used to augment the agent’s context during evaluation, enabling more transition-aware decision making.
LLM-based agents often fail in novel environments due to a semantic mismatch: they lack a “world model” to predict the outcomes of their actions (Ha and Schmidhuber, 2018; Chae et al., 2025). To address this, we propose dynamics grounding, a strategy that proactively discovers the causal dynamics of an environment before the agent encounters any user tasks during deployment. The objective is to construct an in-context world model using natural language by performing a one-time, automated exploration. The cost of this discovery phase is amortized over all subsequent tasks performed in that environment.
@@ 3.2 Dynamics Grounding
    Persona/Exploratory Goal Synthesis. An LLM is prompted with a high-level description of the environment (e.g., website purpose, API documentation) to generate a set of $N$ diverse, exploratory “personas” or “goals”. Personas frame open-ended tasks designed to probe for non-obvious interactions. For our travel-example.com scenario, a synthesized persona might be: “As a first-time user, I want to see what happens if I try to search for a flight without selecting a date.”. We prefer personas over a less-structured exploration policy (e.g. maximizing state novelty or coverage) to guide the agent to explore complex, multi-step interactions that a naive exploration strategy might miss, thus yielding more semantically rich dynamics.
@@
    Exploration and on-the-fly Dynamics Extraction For each persona or goal, an LLM agent interacts with the environment. The agent is instructed to take novel actions to maximize its discovery of new state transitions. After each action is executed and the system transitions into the next state, we take (observation, action, new observation) to summarize into a concise human-readable rule $e$. Immediately after, the list of generated rules $\{e\}_{i=1}^{T}$ is appended to the agent to encourage taking actions that have not been taken before. In our example, the agent would execute click(‘‘Go’’) and log the transition from the main page to a state where a calendar pop-up is active. Then the transition on travel-example.com would be extracted as the dynamic: Executing the action click(‘Go’) on the main page causes a calendar modal to appear for date selection.
@@
    Filtering and Consolidation. The set of all extracted dynamics, $E=\{e\}_{i=1}^{M}$, is passed to a reasoning model. This model filters out trivial (e.g., “typing in a text field shows text”) or repetitive rules, producing a final, clean set of dynamics, $E_{\text{clean}}$.